Alice purchases a boat. The purchase price is $17,300 and she makes a down payment of $3,200. She finances the balance for six years. There are level end-of-month payments (except for a slightly reduced final payment) and the loan is made at an annual effective rate of 5.1%.

(a)

Find the amount of each of the first seventy-one payments. (Round your answer up to the nearest cent.)

(b)

Find the outstanding loan balance just after the twenty-fourth payment. (Round your answer to the nearest cent.)
Answer #1

EAR = 5.1%

So,

APR(monthly) = 12[(1.051)1/12 - 1] = 4.98%

a.

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 14,100, FV = 0, N = 72, I = 0.0498/12]

PMT = $226.95

b.

Calculating Loan Value after 24th Payment,

Using TVM Calculation,

FV = [PV = 14,100, N 24, I = 0.0498/12, PMT = -226.95]

FV = $9,858.63